# Non-Abelian Statistics for Bosonic Symmetry-Protected Topological Phases

### v3

@@ III Braiding statistics of zero modes
In this section, we study the braiding statistics of topological zero modes in the one-dimensional bosonic SPT phase introduced above. Based on the symmetry analysis, we derive the effective Hamiltonian of the braiding process. A central prediction is that the reality condition of Hamiltonian reveals two distinct classes of braiding statistics. We propose two types of tri-junction-based braiding schemes to realize both classes of statistics. The first scheme involves a uniformly half-filled tri-junction, while the second introduces a controllable local defect with empty filling at the tri-junction crossover. Our findings apply to a wide range of bosonic SPT phases with real Hamiltonian realizations.
@@ III.1 General theory of the braiding statistics
Braiding statistics is ill-defined in a simple one-dimensional chain. To address this, we employ a tri-junction configuration, which allows for the controlled braiding of zero modes. The tri-junction consists of two or three spin chains connected at a point, effectively providing an extra spatial dimension that prevents quasiparticle collisions during the braiding process. Details of the tri-junction configuration are presented in the next subsection. The braiding of zero modes is described by the time evolution operator:
$U(T)=\hat{T}e^{-i\int_{0}^{T}H(t)dt},$
where $\hat{T}$ is the time-ordering operator, and $H(t)$ is the time-dependent tri-junction Hamiltonian. The effective Hamiltonian of the braiding process is defined as
@@
Given that $H(t)$ preserves unitary symmetries for any $t$, one can easily verify that the effective Hamiltonian $H_{E}$ must obey the same unitary symmetries. In the case of the topological spin-exchange model, we have $H_{E},U_{\theta}]=[H_{E},C]=0$. Then, $H_{E}$ must take the generic form
@@
$H_{E}=\lambda\left(\beta_{1}^{\dagger}\beta_{2}+\beta_{2}^{\dagger}\beta_{1}\right)+\delta\left(n_{1}+n_{2}-2n_{1}n_{2}\right).$ (15)
Here, we carry over the notation $\beta^{\dagger}_{1}$ ($\beta^{\dagger}_{2}$) from the chain case to denote the zero-mode creation operator localized at the left (right) end of the horizontal chain of the tri-junction, which generally does not admit a simple analytic form as in Eq. (II.1). The occupation number operator of the left (right) zero mode is $n_{1}=\beta_{1}^{\dagger}\beta_{1}$ ($n_{2}=\beta_{2}^{\dagger}\beta_{2}$).
The ground-state degeneracy of the tri-junction inherits from that of the horizontal chain, which hosts two zero modes at the ends. The Fock-space bases read $\{\left|{n_{1},n_{2}}\right\rangle\}$ with $n_{1},n_{2}=0$ or $1$. This degeneracy is topologically protected under symmetry: no local operator can transform these basis states into one another, and no local measurement can fully distinguish them. Local operators that mix or measure the states include $\beta^{\dagger}_{i}$, $\beta_{i}$, $n_{1}$, and $n_{2}$, which however violate the particle-number conservation or particle-hole symmetry of the present bosonic SPT phase. Any symmetry-preserving operations on the ground states must be nonlocal. Therefore, the four-dimensional ground-state subspace is protected by both symmetry and topology, manifesting a symmetry-protected topological subspace. Within this subspace, the symmetry imposes a selection rule analogous to those in topological quantum field theories (TQFTs), where matrix elements of all local operators vanish. This leads to the robustness against local perturbations similar to the topological subspaces in TQFTs. Consequently, when the zero modes are well separated, transitions between different states can only occur through nonlocal braiding operations. Our setup differs fundamentally from conventional TQFT-based frameworks by employing symmetry as a universal mechanism to enforce topological selection rules, thereby extending the traditional framework of non-Abelian statistics.
The parameters $\lambda$ and $\delta$ in Eq. (15) are real to ensure the Hermiticity of $H_{E}$ and are determined by solving the Yang-Baxter equation, subject to the Hamiltonian’s reality condition. As shown in Appendix B, there are two distinct classes of solutions for the effective Hamiltonian. For the first class, the second-quantized form of the braiding operator is
$U^{\text{conv}}=1+\beta_{1}^{\dagger}\beta_{2}+\beta_{2}^{\dagger}\beta_{1}-n_{1}-n_{2}+2n_{1}n_{2}.$ (16)
@@
where the superscript “conv” denotes “conventional”, as will be clarified below. This operator transforms the zero modes as
$U^{\text{conv}}\beta_{1}(U^{\text{conv}})^{\dagger}=\beta_{2},\quad U^{\text{conv}}\beta_{2}(U^{\text{conv}})^{\dagger}=\beta_{1},$ (17)
indicating an exchange of the two zero modes without any additional phase factor. Hence, Eq. (17) transforms the states as $U^{\text{conv}}|0,1\rangle=|1,0\rangle$ and $U^{\text{conv}}|1,0\rangle=|0,1\rangle$, with the full braiding operation satisfying $(U^{\text{conv}})^{2}=1$. Unlike Majorana fermions, no minus sign appears because the zero-mode operators obey hard-core bosonic commutation relations.
A second class of exotic unconventional braiding transformations emerges with the corresponding braiding operator (see Appendix. B):
$U^{\text{def}}$ $=$ $1-\beta_{1}^{\dagger}\beta_{2}+\beta_{2}^{\dagger}\beta_{1}-n_{1}-n_{2}+2n_{1}n_{2}$ (18)
@@
where the superscript “def” denotes “defect-assisted” as to be clarified in the next subsection. Under this braiding operation, the zero modes transform as
@@
and the full braiding recovers a $\pi$ phase factor for each zero mode $(U^{{}^{\text{def}}})^{2}\beta_{1(2)}((U^{\text{def}})^{\dagger})^{2}=-\beta_{1(2)}$. This braiding operation is a fractionalization of the conventional braiding statistics in Eq. (17), exhibiting emergent fermionic zero-mode behavior. Notably, the action of $U^{\text{def}}$ on $\beta_{1}$ and $\beta_{2}$ is nonlinear. Such fractionalized non-Abelian statistics and nonlinear transformations are hallmark features of strongly correlated systems, similar to those hosting parafermion zero modes Hong et al. (2024). This structure can be elucidated via a Jordan-Wigner transformation:
@@
which maps the hard-core bosonic zero modes to Majorana pairs $\gamma_{1}$, $\gamma_{2}$, $\tilde{\gamma}_{1}$ and $\tilde{\gamma}_{2}$. In this representation, the braiding operator becomes
$U^{\text{def}}=\exp\left(-\frac{\pi}{4}\tilde{\gamma}_{1}\tilde{\gamma}_{2}\right)\exp\left(-\frac{\pi}{4}\gamma_{1}\gamma_{2}\right),$ (21)
coinciding with the braiding operator for two pairs of Majorana zero modes in symmetry-protected topological superconductors Liu et al. (2014); Gao et al. (2016); Hong et al. (2022); Haim and Oreg (2019b); Masaki et al. (2024), up to a minus sign in the first exponential. Hence, the second class of fractionalized braiding with $(U^{\text{def}})^{2}=-1$ is analogous to the braiding of Majorana pairs. Whereas Majorana modes transform linearly under braiding, the transformation in Eq. (19)-related through the non-local Jordan-Wigner mapping-contains nonlinear terms.
@@
The nonlinear terms are physically essential by noting that the transformation in Eq. (19) for operators is applicable to the whole four-dimensional ground-state subspace. Owing to particle-number conservation, this subspace separates into three sectors characterized by the total zero-mode occupation number $n_{1}+n_{2}$. The braiding operators act nontrivially in the half-filled subspace sector ($n_{1}+n_{2}=1$) that $U^{\text{def}}|1,0\rangle=|0,1\rangle$ and $U^{\text{def}}|0,1\rangle=-|1,0\rangle$, whereas the $n_{1}+n_{2}=0$ and $2$ sectors are invariant under the braiding transformation, which is ensured by the nonlinear terms in Eq. (19) (see also Appendix B). The unconventional braiding transformation on the four ground states is the same as that for qubit states formed by two Majorana pairs Liu et al. (2014); Hong et al. (2022).
Although the zero-mode operators obey the commutation relation of hard-core bosons, the way that information is stored and manipulated via braiding is fundamentally different from that of local hard-core bosons. This distinction is always overlooked, and we clarify it here. First, the quasiparticles defined by zero edge modes in the topological phase exist regardless of the occupation of the edge modes. This allows us to study braiding in the states $\left|{1,0}\right\rangle$ and $\left|{0,1}\right\rangle$, where braiding a zero-mode quasiparticle in the unoccupied state with one in the occupied state yields non-trivial results. In contrast, a local hard-core boson exists only when the state is occupied, so braiding in the half-filled subspace, which has only one hard-core boson, cannot be defined. Further, as pointed out former, the particle-hole symmetry and particle-number conservation protect that the information encoded in the ground states cannot be affected by any local symmetry-preserving operations. Specifically, for qubit defined from the half-filled sector spanned by $|1,0\rangle$ and $|0,1\rangle$, the information is encoded nonlocally across two zero-mode quasiparticles that are spatially separated by a gapped bulk. The qubit is not locally distinguishable and remains robust against any local symmetry-preserving perturbations, and is protected by the bulk topology of the underlying bosonic SPT phase. Any operator connecting the two basis states, e.g. $\beta_{1}^{\dagger}\beta_{2}+\beta_{2}^{\dagger}\beta_{1}$, is inherently nonlocal. Finally, the two-fold degeneracy of the half-filled sector is also consistent with the equivalence in Eq. (21), in which the two MZM pairs render a two-fold degeneracy in the odd-parity subspace. Therefore, we realize a topologically protected degenerate subspace by encoding states into spatially separated, symmetry- and gap-protected zero-mode occupation. This gives the non-Abelian statistics that cannot be affected by local symmetry-preserving perturbations.
Finally, we show that SPNA statistics can be generalized to a broad class of bosonic SPT phases. From the derivation, it is clear that the effective braiding Hamiltonian depends solely on the symmetries of the system and the reality of the microscopic Hamiltonian. In particular, the latter imposes a crucial constraint, $U^{2}=\pm 1$. Consequently, the two distinct classes of braiding statistics emerge as universal features of bosonic SPT phases described by real Hamiltonians. This universality implies that the observed braiding behaviors and phase differences between ground states are expected to persist across a wide range of systems within this class. Therefore, our results not only apply to the specific model studied but also provide a general framework for understanding SPNA statistics in bosonic SPT phases. The existence of these two distinct classes of braiding statistics constitutes a fundamental and exceptional feature of correlated bosonic SPT phases, distinguishing them from free-fermion topological phases.
@@
Figure 2: Configuration changes of the local defect enclosed by a red dashed line) during the defect-assisted braiding process.
@@ III.2 Braiding schemes with tri-junctions
We propose two types of braiding schemes to realize the two distinct classes of braiding statistics predicted above. For the first braiding scheme, we consider the tri-junction depicted in Fig. 1(c), which consists of three bosonic SPT chains intersecting at a central site. The entire system is uniformly half-filled. For the second braiding scheme, we propose the tri-junction illustrated in Fig. 1(d). The tri-junction comprises a horizontal chain and a vertical chain, both half-filled, along with a local defect. The local defect is controllable in the sense that we can tune its filling to be empty, half-filled, or fully filled. In particular, to realize unconventional statistics, the defect is set to be either empty or fully filled, both yielding identical results due to particle-hole symmetry.
Both tri-junctions, with coupling configurations shown in Figs. 1(c) and 1(d), host two zero modes localized at the horizontal ends of the junction, which induce the four-fold ground-state degeneracy. Among the ground-state manifold, we are particularly interested in the 2D half-filled subspace $\mathcal{H}_{0}=\text{span}\{\left|{1,0}\right\rangle,\left|{0,1}\right\rangle\}$, on which a few remarks are given. First, this 2D subspace respects the particle-hole symmetry and particle-number conservation, making it a natural choice to encode information. Second, as mentioned previously, the subspace $\mathcal{H}_{0}$ can encode a nonlocal qubit under the symmetry constrain. Third, the two braiding operators in Eqs. (16) and (18) in this subspace yield $\sigma_{x}$ and $i\sigma_{y}$ gates, respectively, rendering the non-commutable operations in the single quibt setting. As a consequence, the half-filled subspace $\mathcal{H}_{0}$ is the most physically relevant sector for our study. For simplicity, we henceforth relabel the two basis states as
The states $\left|{\beta_{1}}\right\rangle$ and $\left|{\beta_{2}}\right\rangle$ become exactly orthogonal only in the thermodynamic limit or in the dimerized limit. In the regime $v\gg v^{\prime}$, their overlap $\langle\beta_{1}|\beta_{2}\rangle$ is negligible, so the basis $\left|{\beta_{1}}\right\rangle,\left|{\beta_{2}}\right\rangle$} can be treated as approximately orthonormal. The braiding of zero modes is achieved by adiabatically moving their localization positions, as illustrated in Fig. 1(c) [or Fig. 1(d)]. The braiding process begins by moving the left zero mode to the bottom of the tri-junction, followed by moving the right zero mode to the left end, and finally transferring the bottom zero mode to the right end. This procedure is decomposed into a sequence of elementary steps, each shifting the zero mode by two sites, as depicted in Fig. 1(e). The dynamics of the elementary moving step involving bosonic modes $b_{1,2,3}$ are governed by the Hamiltonian
@@
The braiding results are characterized by the evolution of ground-state wavefunctions $\left|{\beta_{1}(T)}\right\rangle=U(T)\left|{\beta_{1}}\right\rangle$ and $\left|{\beta_{2}(T)}\right\rangle=U(T)\left|{\beta_{2}}\right\rangle$, with $T$ the braiding time. For the first type of braiding scheme, the tri-junction in Fig. 1(c) is a half-filled bosonic SPT system. Due to particle-hole symmetry, the dynamics of ground states $\left|{\beta_{1}}\right\rangle$ and $\left|{\beta_{2}}\right\rangle$ are identical, resulting in no net geometric phase difference between $\left|{\beta_{1}(T)}\right\rangle$ and $\left|{\beta_{2}(T)}\right\rangle$. The ground states $\left|{\beta_{1}}\right\rangle$ and $\left|{\beta_{2}}\right\rangle$ transform as follows after braiding:
@@
We now consider the scheme for the second class of braiding statistics, as illustrated in Fig. 1(d), where the half-filled horizontal chain supports two zero modes at the ends, and the empty local defect at the junction hosts a bosonic quasi-hole. During the braiding, the local defect changes its position around the junction, while remaining isolated from the remaining part of the tri-junction at the termini, and returns to the original state after a single braiding [Fig. 2]. This fulfills the criteria for the well-defined braiding operations. The two zero modes successively pass through the quasi-hole hosted by the local defect. One zero mode goes over once, whereas the other goes over twice. This process induces a $\pi$-phase difference between $\left|{\beta_{1}(T)}\right\rangle$ and $\left|{\beta_{2}(T)}\right\rangle$. The ground states $\left|{\beta_{1}}\right\rangle$ and $\left|{\beta_{2}}\right\rangle$ finally transform as
@@
up to an overall dynamical phase. The minus sign in the above transformation implies an emergent fermionic type of statistics, which aligns with the operator transformation in Eq. (19). We refer to this novel braiding scheme as defect-assisted braiding. It is convenient to represent the braiding results as matrices in the following basis:
@@
According to the analysis above, in the adiabatic limit ($T_{0}\rightarrow+\infty$), the braiding of zero modes can give rise to two possible results (up to a global dynamical phase):

### v4

@@ III Braiding statistics of zero modes
In this section, we study the braiding statistics of topological zero modes in the one-dimensional bosonic SPT phase introduced above. Based on the symmetry analysis, we derive the effective Hamiltonian of the braiding process. A central prediction is that the reality condition of Hamiltonian reveals two distinct classes of braiding statistics. We propose two types of tri-junction-based braiding schemes to realize both classes of statistics. The first scheme involves a uniformly half-filled tri-junction, while the second introduces a controllable local defect with empty filling at the tri-junction crossover. Our findings apply to a wide range of bosonic SPT phases with real Hamiltonian realizations
@@ III.1 General theory of the braiding statistics
Braiding statistics cannot be implemented in a single one-dimensional chain. To address this, we employ a tri-junction configuration, which allows for the controlled braiding of zero modes. The tri-junction consists of two or three spin chains connected at a point, effectively providing an extra spatial dimension that prevents quasiparticle collisions during the braiding process. Details of the tri-junction configuration are presented in the next subsection. The braiding is an unitary evolution which can be characterized by the evolution operator
$U(T)=\hat{T}e^{-i\int_{0}^{T}H(t)dt},$ (12)
where $\hat{T}$ denotes the time-ordered integral for the Hamiltonian $H(t)$ of the tri-junction, and $T$ is the total braiding time. The braiding evolution can further be characterized through an effective Hamiltonian defined as
@@
One can easily verify that the unitary symmetries of $H(t)$ at any $t$ are also preserved for the effective Hamiltonian $H_{E}$. These symmetries constrain the form of $H_{E}$. Furthermore, the braiding operators for different zero modes must satisfy the Yang–Baxter equation
$U_{12}U_{23}U_{12}=U_{23}U_{12}U_{23}.$ (14)
where $U_{i(i+1)}$ denotes the counterclockwise braiding transformation between the $i$th and $(i+1)$th zero modes. For the spin-exchange model, we write down the most generic form of the effective Hamiltonian
@@
$H_{E}=\lambda\bigr(e^{i\alpha}\beta_{1}^{\dagger}\beta_{2}+h.c.\bigr)+\delta\left(n_{1}+n_{2}-2n_{1}n_{2}\right),$ (15)
where the real parameters $\lambda$, $\delta$, and $\alpha$ will be determined by the symmetry constraint and generic fundamental features of the braiding process, and $\beta^{\dagger}_{1}$ ($\beta^{\dagger}_{2}$) denotes the left (right) zero-mode creation operator, with the particle number operator defined by $n_{1}=\beta_{1}^{\dagger}\beta_{1}$ ($n_{2}=\beta_{2}^{\dagger}\beta_{2}$). Substituting $H_{E}$ into the Yang-Baxter equation, we obtain the general form of braiding operator after some algebra (see Methods. VIII.2 for details):
$U=I$ $+$ $ie^{-i\delta T}\left(e^{i\alpha}\beta_{1}^{\dagger}\beta_{2}+e^{-i\alpha}\beta_{2}^{\dagger}\beta_{1}\right)$ (16)
- $n_{1}-n_{2}+2n_{1}n_{2}.$
For the real Hamiltonian $H(t)$, the zero modes after braiding must keep to be real. This enforces that under a full braiding, the transformation may generally take two possibilities that $U^{2}=\pm I$, which corresponds to the two possible Berry phases 0 or $\pi$. Under this constraint the parameters $\delta$ and $\alpha$ are quantified into two sets of discrete values, giving the two distinct classes of braiding statistics which we show below in detail. The two classes of braiding statistics emerge as a universal feature of the bosonic SPT phases, and constitute a fundamental dynamical observable for a broad class of bosonic SPT phases with real Hamiltonians.
We return to the present specific spin-exchange model. For the first class with $U^{2}=I$, one has $\delta=\frac{2n\pi}{T}+\frac{\pi}{2T}$ with $n\in\mathbb{Z}$ and $\alpha=0$, and the second-quantized form of the braiding operator is (see Methods. VIII.2)
$U^{\text{conv}}=I+\beta_{1}^{\dagger}\beta_{2}+\beta_{2}^{\dagger}\beta_{1}-n_{1}-n_{2}+2n_{1}n_{2}.$ (17)
@@
Here we use the superscript “conv” to denote “conventional” that will be clarified below. The above braiding transforms the zero modes as
$U^{\text{conv}}\beta_{1}(U^{\text{conv}})^{\dagger}=\beta_{2},\ U^{\text{conv}}\beta_{2}(U^{\text{conv}})^{\dagger}=\beta_{1},$ (18)
which follows the standard bosonic commutation relation of the two bosonic zero modes. This is in contrast to the braiding of MZMs, in which a minus sign appears since the Majorana zero-mode operators obey fermion commutation relation.
The more intriguing prediction is that the bosonic zero modes allow a second class of exotic unconventional braiding transformations, corresponding to $\delta=\frac{2n\pi}{T}$ with $n\in\mathbb{Z}$ and $\alpha=\frac{\pi}{2}$ in Eq. (16), given by the braiding operator (see Methods. VIII.2):
$U^{\text{def}}$ $=$ $I-\beta_{1}^{\dagger}\beta_{2}+\beta_{2}^{\dagger}\beta_{1}-n_{1}-n_{2}+2n_{1}n_{2}$ (19)
@@
Here the superscript “def” denotes “defect-assisted”, as we shall see in the next subsection that the unconventional braiding statistics can be achieved with assistance by a controlled defect. Under this braiding operation, the zero modes transform as
@@
and the full braiding recovers a $\pi$ phase factor for each zero mode $(U^{{}^{\text{def}}})^{2}\beta_{1(2)}((U^{\text{def}})^{\dagger})^{2}=-\beta_{1(2)}$, rendering $U^{2}=-I$. This nonlinear braiding operation is a fractionalization of the conventional braiding statistics in Eq. (20). The nonlinear transformations is also a hallmark feature of fractionalization in correlated systems, similar to those hosting parafermion zero modes Hong et al. (2024). To further show the nontrivial essence of the unconventional braiding, we map $U^{\text{def}}$ to the braiding of two pairs of symmetry-protected MZMs Liu et al. (2014); Gao et al. (2016); Hong et al. (2022); Haim and Oreg (2019b); Masaki et al. (2024). This structure can be elucidated via a Jordan-Wigner transformation on the zero modes
@@
which maps the hard-core bosonic zero modes to Majorana modes $\gamma_{1}$, $\gamma_{2}$, $\tilde{\gamma}_{1}$ and $\tilde{\gamma}_{2}$. We further obtain
$U^{\text{def}}=\exp\bigr(-\frac{\pi}{4}\tilde{\gamma}_{1}\tilde{\gamma}_{2}\bigr)\exp\bigr(-\frac{\pi}{4}\gamma_{1}\gamma_{2}\bigr).$ (22)
This novel equivalence demonstrates that the unconventional braiding manifests the emergent fermionic zero-mode behavior. While the Majorana modes transform linearly under braiding, the transformation in Eq. (20) related to the Majorana braiding through Jordan-Wigner mapping naturally contains nonlinear terms.
@@
The braiding transformations on the ground states can be obtained straightforwardly. Owing to particle-number conservation, the ground-state space separates into three sectors characterized by the total zero-mode occupation number as $n_{1}+n_{2}=0,1,2$. The braiding operators act nontrivially on the half-filled subspace sector ($n_{1}+n_{2}=1$) that $U^{\text{def}}|1,0\rangle=|0,1\rangle$ and $U^{\text{def}}|0,1\rangle=-|1,0\rangle$, while for the conventional braiding we have $U^{\text{conv}}|1,0\rangle=|0,1\rangle$ and $U^{\text{conv}}|0,1\rangle=|1,0\rangle$. In comparison, the two states $|00\rangle$ and $|11\rangle$ are invariant under both braiding transformations. For $U^{\text{def}}$ this invariance is nontrivial and ensured by the nonlinear terms in Eq. (20). We point out that the unconventional braiding transformation $U^{\text{def}}$ on the four degenerate states is the same as the braiding transformation of two Majorana pairs on the four ground states in the symmetry-protected topological superconductors Liu et al. (2014); Hong et al. (2022). This also explains why the unconventional braiding can be mapped to the SPNA statistics of Majorana pairs as shown above.
The existence of two classes of SPNA statistics facilitates to realize non-commuting topological gates in the single-qubit Hilbert space. The 22D half-filled subspace $\mathcal{H}_{0}=\text{span}\{\left|{1,0}\right\rangle,\left|{0,1}\right\rangle\}$ naturally serves as the qubit space, on which two remarks are given. First, as explained below, the subspace $\mathcal{H}_{0}$ encodes a nonlocal qubit under the symmetry protection. Second, the two braiding operators in Eqs. (17) and (19) in this subspace yield $\sigma^{x}$ and $i\sigma^{y}$ gates, respectively, rendering the non-commutable operations in the single quibt setting. For simplicity, we relabel the two basis states as
In comparison, as is well-known for the MZMs, a single qubit can be encoded by four zero modes, and the non-commuting gates are realized by braiding different pairs of the four Majorana modes Ivanov (2001a); Sarma et al. (2005); Alicea et al. (2011). Here in our setting two hard-core bosonic zero modes are sufficient to encode one qubit and realize non-commutable gates.
We conclude this subsection with a few important remarks on the generic theory for the SPNA statistics of bosonic zero modes in the bosonic SPT phases. In general, the non-Abelian statistics of quasiparticles are obtained when the following three conditions are satisfied Nayak et al. (2008). First, the presence of quasiparticles leads to ground state degeneracy, namely, the quantum dimension of the quasiparticles is greater than one. Second, the unitary transformation between two different degenerate ground states cannot be achieved by local operations, but can be achieved by nonlocal braiding of spatially separated quasiparticles. Third, the braiding operations are topological (discrete). It is clear that the conditions I and III are satisfied by the bosonic zero modes in the present SPT phases. The second condition is regarding the nonlocality of the quasiparticle states. We note that, in general, a non-local state is defined as the one that cannot be changed by local operators under certain conditions, but can be transformed by nonlocal operations. In the framework of Majorana modes, the non-locality of the states refers to that such states cannot be changed by the local operations respecting the topological gap. Actually, if the local perturbation is dynamical, for instance, for the local dynamical noise which couples the local Majorana mode to the bulk, the effective nonlocal couplings can be resulted even only local perturbations are introduced, and the topological quibts will not be stable in general MajoranaNoise . For the qubits defined on zero modes in the present bosonic SPT phase, the key constrain is the symmetry protection. Even though each single zero mode is a local state, as long as the unitary symmetry is preserved, the two-mode qubit states, which define the ground state degeneracy, cannot be changed by any local operators. In other words, it is the symmetry protection that enforces that the ground states become nonlocal and cannot be changed by any local transformation, but can be transformed by braiding. As a result, all the three conditions are satisfied by the bosonic zero modes of the present SPT phases, rendering the rigorous theory of the SPNA statistics uncovered here.
The two classes of non-Abelian braiding statistics of the bosonic zero modes also update the basic understanding of the bosonic SPT phases. For the fundamental theory of bosonic SPT phases, research has mainly focused on developing the topological classification and characterization of these phases through the topological invariants, bulk-boundary correspondence, and symmetry protection mechanisms, etc. While the bosonic topological boundary modes have been widely studied, the braiding statistics of such modes were not considered before. The present prediction shows that, due to symmetry protection, the braiding statistics of topological zero modes in bosonic SPT phases are highly nontrivial and generally fall into two distinct classes, one of which is a fractionalization of the other in real-Hamiltonian systems. The presence of the two distinct classes of SPNA statistics is also a fundamental nature of the bosonic zero modes in the correlated systems, providing a fundamental dynamical character of the bosonic SPT phases.
@@
Figure 2: Configuration changes of the local defect (enclosed by a red dashed line) during the defect-assisted braiding process.
@@ III.2 Braiding schemes with tri-junctions
While the generic theory predicts two distinct classes of non-Abelian statistics, devising explicit schemes to realize them remains highly nontrivial. In this subsection we propose two types of braiding schemes to realize the two distinct classes of braiding statistics. In particular, the conventional braiding statistics can be realized with a standard tri-junction depicted in Fig. 1(c), which consists of three bosonic SPT chains intersecting at a central site. The entire system is uniformly half-filled. To realize unconventional braiding statistics, we propose a defect-assisted tri-junction illustrated in Fig. 1(d). The defect is located at the junction point and formed by two sites, connecting the horizontal and vertical topological chains. In contrast to the half-filling condition in the topological chains, the local defect is set to be either empty or fully filled (both yielding identical results). We find that the defect leads to a nontrivial non-Abelian Berry phase for the zero modes transporting through the tri-junction of this correlated system, for which the unconventional braiding statistics is realized.
Both tri-junctions, with initial coupling configurations shown in Figs. 1(c) and 1(d), host two zero modes localized at the horizontal ends of the junction, which induce the ground-state degeneracy. The braiding of zero modes is achieved by adiabatically moving their localization positions, as illustrated in Fig. 1(c) [or Fig. 1(d)]. The braiding process begins by moving the left zero mode to the bottom of the tri-junction, followed by transporting the right zero mode to the left end, and finally the bottom zero mode to the right end. This procedure is decomposed into a sequence of elementary steps, each shifting the zero mode by two sites, as depicted in Fig. 1(e). The dynamics of the elementary moving step involving bosonic modes $b_{1,2,3}$ are governed by the Hamiltonian
@@
For the first type of braiding scheme, the tri-junction in Fig. 1(c) is in the ground state of a half-filled bosonic SPT system. Due to particle-hole symmetry, the dynamics of ground states $\left|{\beta_{1}}\right\rangle$ and $\left|{\beta_{2}}\right\rangle$ are identical, resulting in no net geometric phase difference between $\left|{\beta_{1}(T)}\right\rangle$ and $\left|{\beta_{2}(T)}\right\rangle$. The ground states $\left|{\beta_{1}}\right\rangle$ and $\left|{\beta_{2}}\right\rangle$ transform as follows after braiding:
@@
We now consider the scheme for the second class of braiding statistics, as illustrated in Fig. 1(d), where the half-filled horizontal chain supports two zero modes at the ends, together with an empty local defect at the junction. During the braiding, the local defect changes its position around the junction, while remaining isolated from the remaining part of the tri-junction at the termini, and returns to the original state after a single braiding [Fig. 2]. This fulfills the criteria for the well-defined braiding operations. Braiding the two zero modes in the presence of a local defect induces a $\pi$ phase difference between $\left|{\beta_{1}(T)}\right\rangle$ and $\left|{\beta_{2}(T)}\right\rangle$. This phase is topological in origin, as the defect is located at the geometric singularity of the tri-junction and alters the topology of the braiding process. As shown in Methods VIII.3, we compute the non-Abelian Berry phases in a tri-junction setting. In particular, we show that the two classes of statistics fall into distinct topological classes, the classification of which can be related to the Stiefel–Whitney classes of real vector bundles. The defect at the junction acts as a singular point, and leads to a topological Berry phase $e^{i\Delta\phi}=\langle\beta_{2}|U(T)|\beta_{1}\rangle\langle\beta_{1}|U(T)|\beta_{2}\rangle$, with $\Delta\phi=\pi$ for the unconventional braiding statistics. Supplementary S1 presents an analytic proof of the braiding results for larger size tri-junctions in the dimerized limit, while the topological $\pi$ phase persists away from the dimerized limit. With the topological $\pi$ phase, the ground states $\left|{\beta_{1}}\right\rangle$ and $\left|{\beta_{2}}\right\rangle$ transform as
@@
up to an overall dynamical phase. The minus sign in the above transformation implies an emergent fermionic type of statistics, which aligns with the operator transformation in Eq. (20). We refer to this novel braiding scheme as defect-assisted braiding.
It is convenient to represent the braiding results as matrices in the following basis:
@@
According to the analysis above, in the adiabatic limit ($T_{0}\rightarrow+\infty$), the braiding of zero modes gives rise to two distinct matrices (up to a global dynamical phase):